OpenStudy (anonymous):

The consecutive integers have a sum of 51. find the intergers

OpenStudy (anonymous):

How many integers?

OpenStudy (texaschic101):

we need to know how many integers ?

OpenStudy (anonymous):

3

OpenStudy (texaschic101):

do you know what consecutive means ?

OpenStudy (anonymous):

following continuously

OpenStudy (anonymous):

x+(x+1)+(x+2)=51

OpenStudy (texaschic101):

correct.. so let the 1st integer be x, and the 2nd integer be x + 1, and the third integer be x + 2. together, they equal 51. x + x + 1 + x + 2 = 51 -- combine like terms 3x + 3 = 51 -- subtract 3 from both sides 3x = 51 - 3 3x = 48 -- divide both sides by 3 x = 48/3 x = 16 x + 1 = 16 + 1 = 17 x + 2 = 16 + 2 = 18 check.. 16 + 17 + 18 = 51 51 = 51 (correct) integers are : 16,17,18 any questions ?

OpenStudy (anonymous):

no thanks